OpenStudy (sidmanfu):

If Baycrest Car Dealership sells 5 silver cars for every 2 red ones, how many silver cars did they sell last year if they sold 50 red cars?

OpenStudy (anonymous):

5/2 = x/50 => x = 250/2 = 125

OpenStudy (sidmanfu):

thanks

OpenStudy (anonymous):

Medal? :)

OpenStudy (sidmanfu):

but how did you get that answer??

OpenStudy (anonymous):

Well, it says for every 5 silver cars, 2 red cars are sold. So, the ratio of silver to red must always be 5/2. If 50 red cars were sold, then 5/2 times silver cars must have been sold. 50 * 5/2 = 125

OpenStudy (sidmanfu):

okay thanks